# Orbital Optical Raman Lattice

### v2

@@ Introduction.
Here, the nontrivial interplay between the $p$-orbital interaction in Eq.(2) and the 2D SOC leads to the emergence of exotic orbital orders with distinct novel topology in the condensate state $|G\rangle$. The $p$-orbital interaction in the present orbital optical Raman lattice tends to generate uniform non-zero orbital angular momentum (OAM) order $\langle G|L_{i,z}|G\rangle$ ($L_{i,z}=-\mathrm{i}b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}+\mathrm{H.c.}$) [66, 80] to reduce the interaction energy, leading to $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=\mathrm{\pm i}\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$. However, the Raman-induced SOC in Eq.(1) prefers $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=\pm\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$, suppressing the uniform OAM $\langle L_{i,z}\rangle$. Thus, it is this competitive relationship between them that gives rise to three distinct phases: (1) When the SOC dominates over the $p$-orbital interaction, the condensate will occur at one of the $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) with $\beta_{\tilde{\bm{k}}_{c},p_{x\downarrow}}=\pm\beta_{\tilde{\bm{k}}_{c},p_{y\downarrow}}$, resulting in a $C_{4}$-symmetry-broken superfluid phase without OAM order; (2) In contrast, if the $p$-orbital interaction plays the leading role, the bosons condense at the $\bm{\Gamma}$ point with $\beta_{\bm{\Gamma},p_{x\downarrow}}=\mathrm{-i}\beta_{\bm{\Gamma},p_{y\downarrow}}$, and the condensate exhibits uniform OAM order; (3) Importantly, in the intermediate regime, where the $p$-orbital interaction and SOC are comparably strong, the bosons condense equally at the $\tilde{\bm{k}}_{c}^{n}$ ($n=1$ $\sim$ 4), with the orbital parameters given by $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=-(-1)^{n}\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$, causing a novel OAM density-wave pattern known as the SOSS phase. Fig.2(b) shows the phase diagram plotted versus condensate density $n_{0}$ and $m_{z}$, which govern the $p$-orbital interaction. A most important feature of the phases in (2) and (3) [orange and green areas in Fig.2(b)] is that they host different types of topological quasiparticles, whose emergence necessitates both the $p$-orbital interaction and the Raman induced 2D SOC. We examine them below in detail.
@@
Adjust the Zeeman splitting $m_{z}$ such that onsite energy $\epsilon_{p\downarrow}(\bm{\mathrm{\Gamma}})$ is well below $\epsilon_{s\uparrow}(\bm{\mathrm{\Gamma}})$. With the dominant $p$-orbital population, the $p$-orbital interaction governs the ground state. Consequently, bosons condense at the $\bm{\mathrm{\Gamma}}$ with orbital parameters $\beta_{\bm{\mathrm{\Gamma}},s_{\uparrow}}=0$ and $\beta_{\bm{\mathrm{\Gamma}},p_{x\downarrow}}=\mathrm{-i}\beta_{\bm{\mathrm{\Gamma}},p_{y\downarrow}}$. As shown in Fig.3(a), this ground state exhibits a uniform OAM order $\langle L_{\bm{i},z}\rangle=2n_{0}|\beta_{\bm{\mathrm{\Gamma}},p_{y\downarrow}}|^{2}$ in real space, rendering the UAMSF which breaks time-reversal symmetry. While the ground state is dominated by $p$-wave interaction, the quasiparticle excitations are actually governed by both the $p$-wave interaction and the 2D SOC, and exhibit nontrivial topology characterized by Chern numbers.
@@
Figure 4: (a) Spatial distribution of $\langle L_{\bm{i},z}\rangle$ for 2D SOSS phase. The positive ($\langle L_{\bm{i},z}\rangle>0$) and negative ($\langle L_{\bm{i},z}\rangle<0$) OAM clusters are staggered, with $\sum_{\mathrm{Global}}\langle L_{\bm{i},z}\rangle=0$. Inset: the relative phase difference $\theta_{i}=\text{Im}\Big{[}\mathrm{ln}\langle b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}\rangle\Big{]}$ between the $p_{x}$ and $p_{y}$ orbitals along path $\bf{r}$. (b) Topological edge excitations of the positive A and negative B clusters. The arrows represent the flow directions of the edge currents. (c) Spin field $(\langle\sigma_{\bm{i},x}\rangle,\langle\sigma_{\bm{i},y}\rangle)$ displays a topological pattern (skyrmion). (d) Spin component $\langle\sigma_{\bm{i},z}\rangle$ presents another staggered high-and low-density clusters pattern. Here, parameters are set as $m_{z}=0.06E_{r}$, $n_{0}=2.0$. And the condensation momenta are $\widetilde{\bm{k}}_{c}^{1}=(k_{c},k_{c})$, $\widetilde{\bm{k}}_{c}^{2}=(-k_{c},k_{c})$, $\widetilde{\bm{k}}_{c}^{3}=(-k_{c},-k_{c})$, and $\widetilde{\bm{k}}_{c}^{4}=(k_{c},-k_{c})$. Note that $k_{c}=\pi/10$.
@@
In order to obtain the quasiparticle excitation spectrum, here we employ the Bogoliubov theory to derive the Bogoliubov-de Gennes (BdG) Hamiltonian ^$\hat{\mathcal{H}}_{\mathrm{BdG}}=\frac{1}{2}\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}\Psi_{\bm{k}}$ [80], where $\Psi_{\bm{k}}^{\dagger}=(b_{\bm{\Gamma}+\bm{k},l,\sigma}^{\dagger},b_{\bm{\Gamma}-\bm{k},l,\sigma})$ denotes the Nambu basis. Performing the Bogoliubov transformation, we have $T_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}(\bm{k})T_{\bm{k}}=E_{\bm{k}}$. Here para-unitary matrix $T_{\bm{k}}$ satisfies $T_{\bm{k}}^{\dagger}\tau_{z}T_{\bm{k}}=\tau_{z}$ ($\tau_{z}=\sigma_{z}\otimes\mathrm{\bm{I}}_{3\times 3}$) and the diagonal terms of $E_{\bm{k}}$ represent the excitation spectrum [83]. As shown in Fig.3 (b), an interaction-induced topological gap opens at the $\bm{\Gamma}$ point [81] and, together with the 2D SOC, separates the first band from the second and third bands, leading to the Chern number of the first band $\mathrm{Ch}^{1}=1$, while the total topological number of the second and third bands $\mathrm{Ch}^{2+3}=-1$. Consequently, these topological excitations are chiral.
Two-dimensional spin-orbital supersolid.
@@
In the intermediate regime with moderate $m_{z}$, the $p$-orbital interaction and the Raman-induced SOC are competitive, and lead to an equal condensation at momenta $\widetilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) with the condensation parameters satisfying $|\gamma_{\widetilde{\bm{k}}_{c}^{n}}|^{2}=\frac{1}{4}$ and $\beta_{\widetilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=-(-1)^{n}\beta_{\widetilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$. Here the corresponding OAM order, which is given by $\langle L_{\bm{i},z}\rangle=2n_{0}\sum_{n,n^{\prime}=1}^{4}\mathrm{Im}[e^{\mathrm{-i}(\tilde{\bm{k}}_{c}^{n}-\tilde{\bm{k}}_{c}^{n^{\prime}})\cdot\bm{i}}\gamma_{\tilde{\bm{k}}_{c}^{n}}^{\ast}\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}^{\ast}\gamma_{\tilde{\bm{k}}_{c}^{n^{\prime}}}\beta_{\tilde{\bm{k}}_{c}^{n^{\prime}},p_{y\downarrow}}]$, exhibits a staggered positive and negative clusters distribution shown in Fig.4(a), with the periods of the two types of clusters $\pi/k_{c}$ in $x$ and $y$ directions. The staggered clusters pattern satisfies $\sum_{\mathrm{Cluster}}\langle L_{\bm{i},z}\rangle\neq 0$ but $\sum_{\mathrm{Global}}\langle L_{\bm{i},z}\rangle=0$, which means that the time-reversal symmetry is broken within each cluster but recovered by averaging for the global system. The supersolid phase exhibits nontrivial topology at each OAM cluster, as characterized by the chiral and anti-chiral edge states of topological excitations on the boundaries of the OAM clusters A ($\langle L_{\bm{i},z}\rangle>0$) and B ($\langle L_{\bm{i},z}\rangle<0$), respectively [see Fig.4(b)] [80]. The chiral and anti-chiral edge states are protected by the opposite Chern numbers of the bulk topological excitations in the A and B clusters, and they are robust under impurities or disorders.
@@
The present supersolid phase embodies rich and exotic features beyond the OAM order. In particular, we investigate the real-space spin texture $\langle\bm{\sigma}_{i}\rangle=(\langle\sigma_{i,x}\rangle,\langle\sigma_{i,y}\rangle,\langle\sigma_{i,z}\rangle)$ of the condensate, with the three components given by $(\langle b_{i,s,\uparrow}^{\dagger}\sum_{\nu}b_{i,p_{\nu},\downarrow}+\mathrm{H.c.}\rangle,\langle-\mathrm{i}b_{i,s,\uparrow}^{\dagger}\sum_{\nu}b_{i,p_{\nu},\downarrow}+\mathrm{H.c.}\rangle,\langle n_{i,s,\uparrow}-\sum_{\nu}n_{i,p_{\nu},\downarrow}\rangle)$ with $\nu=x,y$. Here, the spin texture in $x-y$ plane $(\langle\sigma_{i,x}\rangle,\langle\sigma_{i,y}\rangle)$ given in Fig.4(c) shows a topological vortex lattice structure characterized by vortex ($B$) and anti-vortex ($A$) distributions, corresponding to the negative (B) and positive (A) OAM cluster centers, respectively. Further, the spin component along $z$ direction $\langle\sigma_{iz}\rangle$ in Fig.4(d) exhibits another staggered clusters pattern, with a period of $1/\sqrt{2}$ that of OAM patterns. From the three spin components the combined whole spin texture $\langle\bm{\sigma}_{i}\rangle$ forms a topological skyrmion lattice structure whose periodicity matches the OAM pattern. With the nontrivial topology encoded in the real space and in the momentum space (for quasiparticles), the superfluidity of $|G\rangle$ renders an unprecedented topological SOSS phase breaking both lattice translation symmetry and $U(1)$ symmetry, as predicted in the present orbital optical Raman lattice.
@@ Conclusion.
We explore the exotic many-body quantum phases by proposing an orbital optical Raman lattice scheme that innovatively combines the synthetic SOC and $p$-orbital condensates. We find that the competition between the Raman-induced SOC and $p$-orbital interaction can lead to two novel high-orbital condensed states: the UAMSF phase with uniform OAM order, which hosts global chiral edge excitations, and the 2D SOSS phase, characterized by a staggered OAM pattern and a topological skyrmion structure for spin texture, with opposite topological boundary excitations in neighboring OAM clusters. Unlike the double-well lattice for $p$-orbital systems [69, 71, 72, 75, 76, 77, 78], the present orbital optical Raman lattice exhibits full controllability: both the magnitude and the exact form of the Raman-induced $s-p$ coupling can be independently and precisely controlled, leading to the nontrivial competitive interplay. Furthermore, we design a minimal experimental scheme based on cold atom platform, and the lifetime of the high-orbital many-body quantum phases is also discussed in [80]. Our work provides a new perspective on the further study of SOC and high-orbital physics in optical lattices, and will advance the experimental realization of supersolid phases without long-range interactions.
@@ Acknowledgments.
We thank Xin-Chi Zhou, Ting-Fung Jeffrey Poon, and Ye-Bing Zhang for valuable discussions. This work was supported by National Key Research and Development Program of China (No. 2021YFA1400900), the National Natural Science Foundation of China (Grants No. 12104205, No. 12261160368, and No. 11921005), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000) (Z.H.H., K.H.M., B.Z.W., X.J.L.), and by U.S. AFOSR Grant No. FA9550-23-1-0598 (W.V.L.).

### v3

@@ Introduction.
Next, our calculation finds that the interplay between the $p$-orbital interaction in Eq.(2) and the Raman-induced 2D SOC is highly nontrivial and leads to the emergence of exotic orbital orders with distinct novel topology in the condensate state $|G\rangle$. The $p$-orbital interaction in the present orbital optical Raman lattice tends to generate uniform non-zero orbital angular momentum (OAM) order $\langle G|L_{i,z}|G\rangle$ ($L_{i,z}=-\mathrm{i}b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}+\mathrm{H.c.}$) Liu and Wu (2006); SM to reduce the interaction energy, leading to $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=\mathrm{\pm i}\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$. However, the Raman-induced 2D SOC in Eq.(1) prefers $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=\pm\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$, suppressing the uniform OAM $\langle L_{i,z}\rangle$. Thus, it is this competitive relationship between them that gives rise to three distinct phases: (1) When the SOC dominates over the $p$-orbital interaction, the condensate will occur at one of the $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) with $\beta_{\tilde{\bm{k}}_{c},p_{x\downarrow}}=\pm\beta_{\tilde{\bm{k}}_{c},p_{y\downarrow}}$, resulting in a $C_{4}$-symmetry-broken superfluid phase without OAM order; (2) In contrast, if the $p$-orbital interaction plays the leading role, the bosons condense at the $\bm{\Gamma}$ point with $\beta_{\bm{\Gamma},p_{x\downarrow}}=\mathrm{-i}\beta_{\bm{\Gamma},p_{y\downarrow}}$, and the condensate exhibits uniform OAM order; (3) Importantly, in the intermediate regime, where the $p$-orbital interaction and SOC are comparably strong, the bosons condense equally at the $\tilde{\bm{k}}_{c}^{n}$ ($n=1$sim 4), with the orbital parameters given by $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=-(-1)^{n}\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$, rendering the total system a novel OAM density-wave pattern known as the SOSS phase. Fig.2(b) shows the phase diagram plotted versus condensate density $n_{0}$ and $m_{z}$, which govern the $p$-orbital interaction. A most important feature of the phases in (2) and (3) [orange and green areas in Fig.2(b)] is that they host different types of topological quasiparticles, whose emergence necessitates both the $p$-orbital interaction and the 2D SOC. We examine them below in detail.
@@
Adjust the Zeeman splitting $m_{z}$ such that onsite energy $\epsilon_{p\downarrow}(\bm{\mathrm{\Gamma}})$ is well below $\epsilon_{s\uparrow}(\bm{\mathrm{\Gamma}})$. With the dominant $p$-orbital population, the $p$-orbital interaction governs the ground state. Consequently, bosons condense at the $\bm{\mathrm{\Gamma}}$ with orbital parameters $\beta_{\bm{\mathrm{\Gamma}},s_{\uparrow}}=0$ and $\beta_{\bm{\mathrm{\Gamma}},p_{x\downarrow}}=\mathrm{-i}\beta_{\bm{\mathrm{\Gamma}},p_{y\downarrow}}$. As shown in Fig.3(a), this ground state exhibits a uniform OAM order $\langle L_{\bm{i},z}\rangle=2n_{0}|\beta_{\bm{\mathrm{\Gamma}},p_{y\downarrow}}|^{2}$ in real space, rendering the UAMSF which breaks time-reversal symmetry. While the ground state is dominated by $p$-orbital interaction, the quasiparticle excitations are actually governed by both the $p$-orbital interaction and the 2D SOC, and exhibit nontrivial topology characterized by Chern numbers.
@@
Figure 4: (a) Spatial distribution of $\langle L_{\bm{i},z}\rangle$ for 2D SOSS phase. The positive ($\langle L_{\bm{i},z}\rangle>0$) and negative ($\langle L_{\bm{i},z}\rangle<0$) OAM clusters are staggered, with $\sum_{\mathrm{Global}}\langle L_{\bm{i},z}\rangle=0$. Inset: the relative phase difference $\theta_{i}=\text{Im}\Big{[}\mathrm{ln}\langle b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}\rangle\Big{]}$ between the $p_{x}$ and $p_{y}$ orbitals along path $\bf{r}$. (b) Topological edge excitations of the positive A and negative B clusters. The arrows represent the flow directions of the edge currents. (c) Spin field $(\langle\sigma_{\bm{i},x}\rangle,\langle\sigma_{\bm{i},y}\rangle)$ displays a topological pattern (skyrmion). (d) Spin component $\langle\sigma_{\bm{i},z}\rangle$ presents another staggered high-and low-density clusters pattern. Here, parameters are set as $m_{z}=0.06E_{r}$, $n_{0}=2.0$. Besides, the condensation momenta are $\widetilde{\bm{k}}_{c}^{1}=(k_{c},k_{c})$, $\widetilde{\bm{k}}_{c}^{2}=(-k_{c},k_{c})$, $\widetilde{\bm{k}}_{c}^{3}=(-k_{c},-k_{c})$, and $\widetilde{\bm{k}}_{c}^{4}=(k_{c},-k_{c})$, where $k_{c}=\pi/10$.
@@
In order to obtain the quasiparticle excitation spectrum, here we employ the Bogoliubov theory to derive the Bogoliubov-de Gennes (BdG) Hamiltonian ^$\hat{\mathcal{H}}_{\mathrm{BdG}}=\frac{1}{2}\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}\Psi_{\bm{k}}$ SM , where $\Psi_{\bm{k}}^{\dagger}=(b_{\bm{\Gamma}+\bm{k},l,\sigma}^{\dagger},b_{\bm{\Gamma}-\bm{k},l,\sigma})$ denotes the Nambu basis. Performing the Bogoliubov transformation, we have $T_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}(\bm{k})T_{\bm{k}}=E_{\bm{k}}$. Here para-unitary matrix $T_{\bm{k}}$ satisfies $T_{\bm{k}}^{\dagger}\tau_{z}T_{\bm{k}}=\tau_{z}$ ($\tau_{z}=\sigma_{z}\otimes\mathrm{\bm{I}}_{3\times 3}$) and the diagonal terms of $E_{\bm{k}}$ represent the excitation spectrum Shindou et al. (2013). As shown in Fig.3 (b), an interaction-induced topological gap opens at the $\bm{\Gamma}$ point Sun et al. (2012) and, together with the 2D SOC, separates the first band from the second and third bands, leading to the Chern number of the first band $\mathrm{Ch}^{1}=1$, while the total topological number of the second and third bands $\mathrm{Ch}^{2+3}=-1$. Consequently, these topological excitations are chiral.
Two-dimensional topological spin-orbital supersolid.
@@
In the intermediate regime with moderate $m_{z}$, the $p$-orbital interaction and the Raman-induced SOC are competitive, and lead to an equal condensation at momenta $\widetilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) with the condensation parameters satisfying $|\gamma_{\widetilde{\bm{k}}_{c}^{n}}|^{2}=\frac{1}{4}$ and $\beta_{\widetilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=-(-1)^{n}\beta_{\widetilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$. Here the corresponding OAM order, which is given by $\langle L_{\bm{i},z}\rangle=2n_{0}\sum_{n,n^{\prime}=1}^{4}\mathrm{Im}[e^{\mathrm{-i}(\tilde{\bm{k}}_{c}^{n}-\tilde{\bm{k}}_{c}^{n^{\prime}})\cdot\bm{i}}\gamma_{\tilde{\bm{k}}_{c}^{n}}^{\ast}\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}^{\ast}\gamma_{\tilde{\bm{k}}_{c}^{n^{\prime}}}\beta_{\tilde{\bm{k}}_{c}^{n^{\prime}},p_{y\downarrow}}]$, exhibits a staggered positive and negative clusters distribution shown in Fig.4(a), with the periods of $A(B)$-type clusters being $\pi/k_{c}$ in $x$ and $y$ directions. The staggered clusters pattern satisfies $\sum_{\mathrm{Cluster}}\langle L_{\bm{i},z}\rangle\neq 0$ but $\sum_{\mathrm{Global}}\langle L_{\bm{i},z}\rangle=0$, which means that the time-reversal symmetry is broken within each cluster but recovered by averaging for the global system. The supersolid phase exhibits nontrivial topology at each OAM cluster, as characterized by the chiral and anti-chiral edge states of topological excitations on the boundaries of the OAM clusters A ($\langle L_{\bm{i},z}\rangle>0$) and B ($\langle L_{\bm{i},z}\rangle<0$), respectively [see Fig.4(b)] SM . The chiral and anti-chiral edge states are protected by the opposite Chern numbers of the bulk topological excitations in the A and B clusters, and they are robust under impurities or disorders.
@@
The present supersolid phase embodies rich and exotic features beyond the OAM order. In particular, we investigate the real-space spin texture $\langle\bm{\sigma}_{i}\rangle=(\langle\sigma_{i,x}\rangle,\langle\sigma_{i,y}\rangle,\langle\sigma_{i,z}\rangle)$ of the condensate, with the three components given by $(\langle b_{i,s,\uparrow}^{\dagger}\sum_{\nu}b_{i,p_{\nu},\downarrow}+\mathrm{H.c.}\rangle,\langle-\mathrm{i}b_{i,s,\uparrow}^{\dagger}\sum_{\nu}b_{i,p_{\nu},\downarrow}+\mathrm{H.c.}\rangle,\langle n_{i,s,\uparrow}-\sum_{\nu}n_{i,p_{\nu},\downarrow}\rangle)$ with $\nu=x,y$. Here, the spin texture in $x$--$y$ plane $(\langle\sigma_{i,x}\rangle,\langle\sigma_{i,y}\rangle)$ given in Fig.4(c) shows a topological vortex lattice structure characterized by vortex ($B$) and anti-vortex ($A$) distributions, corresponding to the negative (B) and positive (A) OAM cluster centers, respectively. Further, the spin component along $z$ direction $\langle\sigma_{iz}\rangle$ in Fig.4(d) exhibits another staggered clusters pattern, with a period of $1/\sqrt{2}$ that of OAM patterns. From the three spin components the combined whole spin texture $\langle\bm{\sigma}_{i}\rangle$ forms a topological skyrmion lattice structure whose periodicity matches the OAM pattern. With the nontrivial topology encoded in the real space and in the momentum space (for quasiparticles), the superfluidity of $|G\rangle$ renders an unprecedented 2D topological SOSS phase breaking both lattice translation symmetry and $U(1)$ symmetry, as predicted in the present orbital optical Raman lattice.
@@ Conclusion.
[…]
@@ Acknowledgments.
We thank Xin-Chi Zhou, Ting-Fung Jeffrey Poon, and Ye-Bing Zhang for their valuable discussions. This work was supported by National Key Research and Development Program of China (No. 2021YFA1400900), the National Natural Science Foundation of China (Grants No. 12104205, No. 12261160368, and No. 11921005), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000) (Z.H.H., K.H.M., B.Z.W., X.J.L.), and by U.S. AFOSR Grant No. FA9550-23-1-0598 (W.V.L.).